40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16333     Entry time: Wed Sep 15 23:38:32 2021
Author: Koji 
Type: Update 
Category: ALS 
Subject: ALS ASX PZT HV was off -> restored 

It was known that the Y end ALS PZTs are not working. But Anchal reported in the meeting that the X end PZTs are not working too.

We went down to the X arm in the afternoon and checked the status. The HV (KEPCO) was off from the mechanical switch. I don't know this KEPCO has the function to shutdown the switch at the power glitch or not.
But anyway the power switch was engaged. We also saw a large amount of misalignment of the X end green. The alignment was manually adjusted. Anchal was able to reach ~0.4 Green TRX, but no more. He claimed that it was ~0.8.

We tried to tweak the SHG temp from 36.4. We found that the TRX had the (local) maximum of ~0.48 at 37.1 degC. This is the new setpoint right now.

Attachment 1: P_20210915_151333.jpg  4.350 MB  Uploaded Thu Sep 16 00:50:07 2021  | Hide | Hide all
P_20210915_151333.jpg
ELOG V3.1.3-